OpenStudy (anonymous):

21x^2-20x=0 my answers are 0, 20/21 is this correct?

OpenStudy (phi):

21 x^2 - 20x =0 factor out x x (21 x -20) =0 this means either x=0 or 21 x -20 =0 the 2nd equation is 21 x -20=0 add 20 to both sides 21 x = 20 divide both sides by 21 x= 20/21

OpenStudy (phi):

you could replace x with your answers to check if they are correct. x=0 is easy to check: 21*0*0 - 20*0 is 0 -0 or just 0

OpenStudy (anonymous):

so my answers are correct. :) awesome! finally getting a hang of this.

OpenStudy (anonymous):

thank you for your help! it was very appreciated!